# Preference Distillation for Personalized Generative Recommendation

### v1

@@ Appendix B Training without Task-Specific Prompts
Table 5: Performance comparison on sequential recommendation.
@@
PeaPOD 0.0505 0.0400 0.0611 0.0432 0.0588 0.0445 0.0738 0.0493 0.0692 0.0583 0.0787 0.0609
PeaPOD-X 0.0480 0.0381 0.0581 0.0412 0.0565 0.0402 0.0732 0.0456 0.0705 0.0557 0.0835 0.0594
Table 6: Performance comparison on top-n recommendation.
@@
PeaPOD 0.1212 0.2741 0.2007 0.3616 0.2286 0.1097 0.2441 0.1780 0.3260 0.2039 0.0728 0.1618 0.1190 0.2315 0.1411
PeaPOD-X 0.1232 0.2744 0.2017 0.3662 0.2309 0.1130 0.2480 0.1820 0.3304 0.2079 0.0790 0.1795 0.1308 0.2513 0.1536
Table 7: Performance comparison on explanation generation.
@@
PeaPOD 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
PeaPOD-X 0.9561 12.1112 1.7613 9.4255 0.9598 11.3673 1.7317 8.8236 2.5306 15.3696 3.8183 12.6750

### v2

@@ 6.4. Effects of Model Size (RQ4)
Table 6. Performance comparison on sequential recommendation between small and base models.
@@
PeaPOD-PMF (small) 0.0534 0.0424 0.0638 0.0456 0.0718 0.0527 0.0915 0.0591 0.0865 0.0708 0.1018 0.0750
PeaPOD-PMF (base) 0.0562 0.0461 0.0658 0.0488 0.0582 0.0443 0.0721 0.0488 0.0731 0.0589 0.0827 0.0610
Table 7. Performance comparison on top-n recommendation between small and base models.
@@
PeaPOD-PMF (small) 0.1252 0.2708 0.2006 0.3596 0.2287 0.1292 0.2628 0.1978 0.3423 0.2228 0.0728 0.2126 0.1581 0.2872 0.1817
PeaPOD-PMF (base) 0.1280 0.2698 0.2016 0.3588 0.2300 0.1000 0.2268 0.1652 0.3086 0.1914 0.0649 0.1494 0.1079 0.2168 0.1291
Table 8. Performance comparison on explanation generation between small and base models.
@@
PeaPOD-PMF (small) 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
PeaPOD-PMF (base) 1.0548 14.1369 2.0106 11.1413 0.8602 13.2299 1.3422 10.0028 2.5037 14.4358 3.9266 11.8846
6.5. Effects of Multi-head Attention
Figure 5. The effects of different number of heads used for attention on the Beauty dataset. Note that ‘0’ represents multi-layer perceptron (MLP) rather than the number of heads.
To understand the effects of multi-head attention versus alternative implementations, we perform an ablation experiment with a varying number of heads. Specifically, to create ablated versions of the collaborative user prompt $\boldsymbol{p}$, we train PeaPOD-PMF on the Beauty dataset with the number of heads set to {1, 2, 4, 8, 16}. In addition, we also experiment with passing the user embedding through a multi-layer perceptron (MLP). Rather than performing attention and passing the attended values through a linear layer, MLP passes the user embedding directly through a linear layer to project the user embedding as a personalized soft prompt. We include this experiment to observe if encoding collaborative information via attention provides superior performance versus simpler methods. We show the experimental results in Figure 5.
We observe that multi-head attention with 4 heads, denoted as $MHA_{4}$, achieved the best overall performance, particularly in sequential recommendation. $MHA_{4}$ showed a slight improvement over single-head attention (referred to as ‘1’ in Figure 5) in all explanation generation metrics. For top-n recommendation, the performance of $MHA_{4}$ and single-head attention was close to equal. We also observe that all attention-based models significantly outperformed the MLP implementation, showing the advantage of the compositional attention strategy. However, increasing the number of heads beyond 4 led to a decline in performance for all three recommendation tasks. This suggests that while adding more heads improves performance up to a certain point, exceeding this threshold causes the model to overfit.
7. Conclusion
In this paper, we propose PeaPOD, a novel generative recommender model that effectively distills individual and group user preferences into a collaborative user prompt. In particular, we introduce a multi-head attention strategy that uses the query-key-value mechanism to generate a personalized user embedding. Multi-head attention allows the model to capture muliple user-to-user relationships and enables the model to efficiently share collaborative knowledge between similar users. Our experiments confirm the effectiveness of our PeaPOD framework in three common recommendation tasks. Future directions include adding additional information during training, such item metadata, and developing an end-to-end architecture to unify the user embedding initialization and multi-head attention stages.
8. Limitations and Ethical Concerns
One limitation of our approach is that we restrict our study to exclusively using user and item IDs. Thus, we do not incorporate any additional metadata such as user profiles or item descriptions. We find that this may hinder performance on explanation generation, which relies on having a strong understanding of both the user and item to provide a personalized response. We predict that providing additional metadata to the model can improve performance, particularly if it is given in natural language text as the LLM is already pre-trained on such data. We plan to extend PeaPOD to utilize this information in future work.
Another concern is the issue of fairness in generative recommender systems. For example, prior research has shown that LLM-based recommenders skew towards popular items (Yoon et al., 2024), which can cause a disparity in recommendations and lead to a bias-amplification loop. In addition, past research has investigated the issue of fairness in the LLM itself (Salutari et al., 2023). Thus, careful consideration must be made when deploying a generative recommender model to mitigate unfair recommendations.